OpenStudy (anonymous):

19j - 22 = -7j + 30 J = ?

OpenStudy (anonymous):

j=2

OpenStudy (anonymous):

19j-22=-7j+30 19j+7j=30+22 26j=52 j=52/26 j=2

OpenStudy (anonymous):

Yes ! ^